Judy Hays wants to give $3,200 to the building fund at her local church. Assuming Judy can itemize her deductions, how much will this contribution save her on her federal income taxes if she is in the 28 percent marginal tax bracket?
Answer #1

Answer:

Donation for building her local church D=$3200

Tax rate T=28%

Saving on tax=D*T=3200*28%=$896